# Real-World Offline Reinforcement Learning from Vision Language Model Feedback

### v1

@@ V-B1 Setup
The offline dataset we use is collected from a prior human study in Wang et al. [30]. The policy used to collect the dataset is trained using reinforcement learning and policy distillation in simulation and then transferred to the real world. Due to sim2real gaps, the policy is not optimal in the dressing task, and thus the offline dataset is not optimal and contains failure trajectories. There are in total 485 trajectories, which corresponds to 26158 transitions in this offline dataset. As garments do not have a known compact state representation, the logged observation in the dataset is the segmented point cloud of the scene, which contains the garment point cloud, the arm point cloud, and a point that corresponds to the robot end-effector, which we use to train the reward and the policy. We randomly sampled 4000 image pairs to query the VLM for preference labels. The VLM we use for this real-world dressing task is GPT-4o.
We compare to a state-of-the-art behavior cloning baseline that takes point cloud as input: 3D diffusion policy (DP3) [25]. The evaluation metric is the arm dressed ratio, which is the ratio between the dressed distance on the arm and the total length of the arm. We test both methods in two scenarios: the first is dressing the manikin, which holds a “L” shape arm pose. The second is dressing the ViperX 300 S arm in a pose akin to a human stretching their forearm away from their torso. We note that both of these two arms are not present in the training set, which only includes data collected from real people. The shape and morphology of the manikin arm is similar to those of real people, and the shape and morphology of the ViperX 300 S arm is rather distinct from arms of real people. Each dressing trial stops after 60 steps, or when the sensed force on the Sawyer end-effector is larger than 7 Newtons. We test each method with 3 different garments, as shown in Fig. 3. For both the manikin and the ViperX 300 S arm, we run each method on each garment for 5 times and report the mean and standard deviation of the dressed ratios.
@@ VI CONCLUSIONS
In this paper, we propose Offline RL-VLM-F, a new system that enables automatic reward labeling and policy learning from unlabeled, sub-optimal offline datasets. We build on prior work RL-VLM-F to automatically generate reward labels for offline datasets using preference feedback from a vision-language model and a text description of the task. Our method then learns a policy using offline RL with the reward-labeled dataset. We test our system in a complex real-world robot-assisted dressing task, where we first learn a reward function using a vision-language model on a sub-optimal offline dataset, and then we use the learned reward to employ Implicit Q Learning to develop an effective dressing policy. Our method also performs well in simulation tasks involving the manipulation of rigid and deformable objects, outperforming baselines such as behavior cloning and inverse RL. In summary, we propose a new system that enables automatic reward labeling and policy learning from unlabeled, sub-optimal offline datasets.

### v2

@@ V-B1 Setup
The offline dataset we use is collected from a prior human study in Wang et al. [30]. The policy used to collect the dataset is trained using reinforcement learning and policy distillation in simulation and then transferred to the real world. Due to sim2real gaps, the policy is not optimal in the dressing task, and thus, the offline dataset is not optimal and contains failure trajectories. There are in total 485 trajectories, which correspond to 26158 transitions in this offline dataset. As garments do not have a known compact state representation, the logged observation in the dataset is the segmented point cloud of the scene, which contains the garment point cloud, the arm point cloud, and a point that corresponds to the robot end-effector, which we use to train the reward and the policy. We randomly sampled 4000 image pairs to query the VLM for preference labels. The VLM we use for this real-world dressing task is GPT-4o.
We compare to a state-of-the-art behavior cloning baseline that takes point cloud as input: 3D diffusion policy (DP3) [25]. The evaluation metric is the arm dressed ratio, which is the ratio between the dressed distance on the arm and the total length of the arm. We test both methods in two scenarios: the first is dressing the manikin, which holds a “L” shape arm pose. The second is dressing the ViperX 300 S arm in a pose akin to a human stretching their forearm away from their torso. We note that neither of these two arms is present in the training set, which only includes data collected from real people. The shape and morphology of the manikin arm are similar to those of real people, and the shape and morphology of the ViperX 300 S arm are rather distinct from arms of real people. Each dressing trial stops after 60 steps, or when the sensed force on the Sawyer end-effector is larger than 7 Newtons. We test each method with 3 different garments, as shown in Fig. 3. For both the manikin and the ViperX 300 S arm, we run each method on each garment 5 times and report the mean and standard deviation of the dressed ratios.
@@ V-B2 Real-world results
V-B3 VLM Limitations and Failure Modes
We observed multiple failure modes of our approach. First, the VLM can misinterpret visual context due to camera angles—for example, it occasionally fails to detect the robot grasping the garment and incorrectly assumes the person is already dressed. Second, in early time steps of the dressing trajectories, the visual cues for progress (such as changes in cloth height or deformation) can be too subtle, leading the VLM to produce inaccurate preference labels. Third, we observed hallucinations in certain cases—e.g., the VLM mistakenly identified a hanger in the scene, which was not present. Lastly, when the two queried images are visually similar—especially when sampled from adjacent time steps—the VLM produces noisy preferences, and the query is not informative. To mitigate this, we only compare frames that are beyond a threshold number of timesteps apart. This ensures greater visual distinction between comparisons and improves label quality.
@@ VI CONCLUSIONS
In this paper, we propose Offline RL-VLM-F, a new system that enables automatic reward labeling and policy learning from unlabeled, sub-optimal offline datasets. We build on prior work, RL-VLM-F, to automatically generate reward labels for offline datasets using preference feedback from a vision-language model and a text description of the task. Our method then learns a policy using offline RL with the reward-labeled dataset. We test our system in a complex real-world robot-assisted dressing task, where we first learn a reward function using a vision-language model on a sub-optimal offline dataset, and then we use the learned reward to employ Implicit Q Learning to develop an effective dressing policy. Our method also performs well in simulation tasks involving the manipulation of rigid and deformable objects, outperforming baselines such as behavior cloning and inverse RL. In summary, we propose a new system that enables automatic reward labeling and policy learning from unlabeled, sub-optimal offline datasets.
VII Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. IIS-2046491 and the National Institute of Standards and Technology (NIST) under Grant No. 70NANB24H314. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation or NIST.